OpenStudy (anonymous):

x ( is the day tickets go on sale). Tickets= -0.2x^2 +12x+11 Use the quadratic equation to determine the last day that tickets will be sold. Answer should be in terms of the number of days after ticket sales begin.

OpenStudy (anonymous):

Mid 61st Day You find it by solving for zero and then taking the larger number.

OpenStudy (anonymous):

Is this the right way to show the 61st day? -0.2x^2+12x+11=0 x=[-12+-sqrt(144-4*-0.2*11)]/(-0.4) x=[-12+-sqrt(152.8)]/(-0.4) Then we do x= [-12-12.36]/-0.4 x= 60.9

OpenStudy (anonymous):

The answer is right, I didn't look into your math all that much....

OpenStudy (anonymous):

I believe the answer is right, just wanted to try and figure it out lol Thanks for your help though :)